OpenStudy (anonymous):

2. As you are leaving the community, you notice anohter new communtity just down the street. Because you are in the area, you decide to inquire about it. a. x+y=38 b. 175,000x+200,000y=7,200,00 c. Use elimination to determine how many houses are available in each floor plan. Explain how you arrived at your answer. Is is this right: sixteen with floor plan one (there are x of them and x is 16) and therefore twenty-two with floorplan 2 PLZZZZZ HELP!!!!!

OpenStudy (anonymous):

we did this yes?

OpenStudy (anonymous):

you need the method, right?

OpenStudy (anonymous):

first of all lets make life easy. \[x+y=38\] is fine as it is but \[175000x+200000y=7200000\] is silly. lets change the last one to \[175x+200y=7200\] i just lopped of 3 zeros

OpenStudy (anonymous):

KINDA NEED AN ANSWER!!!!!

OpenStudy (anonymous):

answer is \[x=16, y=22\] you have it written above

OpenStudy (anonymous):

No, what you need is to understand what you are doing so that you can actually solve problems like these on your exam.

OpenStudy (anonymous):

you wrote the answer in the question. i think the real question is how do you get the answer.

OpenStudy (anonymous):

what identity said

OpenStudy (anonymous):

YEAH YEAH YEAH :p THX

OpenStudy (anonymous):

ima hacker not a math leet

OpenStudy (anonymous):

I'm sure you're a real pro haxxor. But as it stands, you aren't capable of solving this equation and since you're here asking this question, it's probably homework. You'll need to do homework more often than this one time, probably. You'll probably even have an exam as well.

OpenStudy (anonymous):

Oh, and by the way: being able to solve a system of 2 equations does in no way make you a "math leet".

OpenStudy (anonymous):

me, i am a math hole

OpenStudy (anonymous):

yeah well i blow at some math

OpenStudy (anonymous):

Good 4 u satellite